A test for the presence in the blood of antibodies to HIV, the virus that causes AIDS, gives a positive result with probability about 0.004 when a person who is free of HIV antibodies is tested. This is called a false positive. A clinic tests 5000 people who are all free of HIV antibodies. What is the probability that between 10 and 15 people in this sample will receive a false positive reading?
Answer #1

olution Giren data n = 5000 P=0.004./qa-p = 1-0.004 -0.996 probabiuity that between lo and i5 people. - Plio sx515) = P(x-10)